OpenStudy (anonymous):

Hae friends, will you show that the function y=x^2 is not one to one?

OpenStudy (callisto):

When x = 1, y = ...?

OpenStudy (callisto):

When x = -1, y = ...?

OpenStudy (anonymous):

To show something is one-to-one, you would need to show that f(a) = f(b) if and only if a = b. So basically we're just showing that this is not true. f(a): (a)^2 f(b): (b)^2 a^2 = b^2 square root both sides and you get \[a = \pm b\], which shows the function is not one-to-one.

OpenStudy (anonymous):

When x=1,y =1

OpenStudy (anonymous):

when x=-1, y=1

OpenStudy (callisto):

See! You have two different inputs generating the same output!

OpenStudy (anonymous):

yes!

OpenStudy (callisto):

By definition, it is not one to one.

OpenStudy (isaiah.feynman):

Graphically x^2 fails the horizontal line test. The graph of a one to one MUST only intersect the horizontal line once and only once!

OpenStudy (anonymous):

thanks guys!